# Distractor-free Generalizable 3D Gaussian Splatting

### v2

@@ 4.1.2 Mask Refinement
Finally, we decouple $\mathcal{M}_{D}$ from $\mathcal{M}_{Q}$ and recombine them after introducing the segmentation model qi2022high and refining the distractor mask. The final refined mask, termed $\mathcal{M}$ in Fig. 3, substitutes $\mathcal{M}_{Rob}$ in Eq. 3 to mitigate distractor effects during training. Furthermore, we observe that if only Mask Refinement is used, without leveraging reference and our observation, the misclassification remains severe in Fig. 3, which verifies the importance of our reference filter in Sec. 4.1.1.
@@
$\mathcal{L}_{A}=\sum_{N}^{i=1}\mathcal{W}_{T\rightarrow i}(1-\mathcal{{M}})\odot\mathcal{M}^{En}_{Ref_{i}}\odot\left\|\mathbf{I}_{i}-\mathcal{G}\left(\mathbf{P}_{i}\right)\right\|_{2}^{2}.$ (8)
@@ 4.2 Distractor-free Generalizable Inference
Despite improvements in training robustness and mask prediction, DGGS’s inference faces two key limitations: (1) insufficient references compromise reliable reconstruction of occluded and unseen regions; (2) persistent distractor in references inevitably appear as artifacts in synthesized novel views for feedforward inference paradigm (encoder-decoder) in Eq. 2. To address these challenges, we propose a two-stage Distractor-free Generalizable Inference framework in Fig. 4. The first stage employs a Reference Scoring mechanism in Sec.4.2.1 to score candidate references from the images pool, facilitating the selected references with minimal distractor and disparity. The second stage uses a Distractor Pruning module in Sec.4.2.2 to suppress the remaining artifacts.
4.2.1 Reference Scoring Mechanism
@@
The objective of the first inference stage is to select references with minimal distractor and disparity among the predefined scene-images pool, adjacent $K$ views ($K\geq N$) sampled in the test scenes. Therefore, we propose a Reference Scoring mechanism based on the pre-trained DGGS in Sec. 4.1. Specifically, it first samples $N$ adjacent references near the query from the scene-images pool $\{\mathbf{I}\}^{K}_{P}$ for coarse 3DGS prediction. We then designate unselected views from $\{\mathbf{I}\}^{K}_{P}$ as query for mask prediction {}$\{\mathcal{M}\}^{K-N}$, while the distractor masks of the chosen reference views are $\{\mathcal{M}^{En}_{Ref}\}^{N}$. All image masks in pool are scored by Eq. 10, in which top $N$ images are selected in next stage,
$\{\mathbf{I}_{i}\}^{N}=\{\mathbf{I}_{i}\}^{K}_{P}\mid i\in\arg\max_{N}\left\{\mathcal{S}\left(\{\mathcal{M}\}^{K-N};\{\mathcal{M}^{En}_{Ref}\}^{N}\right)\right\}.$ (10)
@@
where $\mathcal{S}$ is the pixel-wise summation for each mask. In practice, besides distractor size, the extrinsics of candidate images are also crucial reference scoring factors due to disparity. Thanks to the discussion of disparity-induced error masks in Sec. 4.1.2, we can directly utilize the count of positive pixels in $\mathcal{M}$ as the single criterion, which selects references that provide better coverage of query view, as shown in Fig. 7 10. In the second stage, we employ top-ranked images as references for fine reconstruction, effectively reselecting the candidate images in the pool without increasing GPU memory. Although this approach successfully handles distractor-heavy references, it comes at the cost of decreased rendering efficiency. Optionally, we mitigate this by halving the image resolution in the first stage.
4.2.2 Distractor Pruning
@@
Although cleaner references are selected, obtaining $N$ distractor-free images in the wild is virtually impossible. These residual distractor propagate via the gaussian encoding-decoding process in Eq. 2, manifesting as phantom splats in rendered query view, as shown in Fig. 10. Therefore, we propose a Distractor Pruning protocol in the second inference stage, which is readily implementable given the reference distractor masks in Sec. 4.2.1. Instead of direct masking on the references, which affects one-to-one mapping between pixels and gaussian primitives in Eq. 2, we selectively prune gaussian primitives within the 3D space by directly removing decoded attributes in distractor regions while preserving the remaining components. However, when references exhibit a large amount of commonly occluded regions, the pruning strategy induces white speckle artifacts. Consequently, based on projected masks, DGGS implements the pruning strategy only in scenarios where it is not considered a common occluded region for all references, which is also discussed in Limitation.
@@ 5 Experiments
This section presents qualitative and quantitative results for DGGS in real-world and our synthetic generalization scenarios with distractor, which validate the reliability of our training and inference paradigm. Also, multi-scene experiments show that DGGS enables existing distractor-free methods to generalize in reconstruction and masking, overcoming their lack of cross-scene inference.
@@
Pixelsplat* (2024 CVPR) charatan2024pixelsplat 18.65 0.673 0.254 17.98 0.557 0.364 20.10 0.704 0.279 Pre-Train
Mvsplat* (2024 ECCV) chen2024mvsplat 18.88 0.670 0.225 18.24 0.586 0.301 20.03 0.722 0.255 on Re10K
Pixelsplat (2024 CVPR) charatan2024pixelsplat 15.49 0.378 0.531 16.34 0.331 0.492 16.02 0.422 0.511
Mvsplat (2024 ECCV) chen2024mvsplat 15.05 0.412 0.391 16.17 0.509 0.381 15.45 0.515 0.426
+RobustNeRF (2023 CVPR) sabour2023robustnerf 16.17 0.463 0.382 16.46 0.470 0.411 17.11 0.534 0.400 Re-Train on
+On-the-go (2024 CVPR) Ren2024NeRF 14.73 0.366 0.522 15.05 0.440 0.472 15.44 0.476 0.526 Distractor-
+NeRF-HuGS (2024 CVPR) chen2024nerf 18.21 0.694 0.266 18.33 0.640 0.299 19.18 0.700 0.283 Datasets
+SLS (Arxiv 2024) sabour2024spotlesssplats 18.11 0.695 0.270 18.84 0.662 0.282 19.29 0.709 0.286
@@
$+$Mask Refinement (DGGS-TR) 21.02 0.738 0.242
@@
Table 4: Ablations for Reference Scoring.
Methods Mean (RobustNeRF)
[…]
DGGS-TR ($N$=4) 21.02 0.738 0.242
DGGS-TR ($N$=8) 20.22 0.690 0.311
DGGS w/o Pruning ($N$=4) 21.47 0.749 0.242
@@
5.1 Experimental Details
5.1.1 Real and Synthetic Distractor Datasets
In accordance with existing generalization frameworks, DGGS is trained in a cross-scene setting with distractor and evaluated in novel unseen distractor scenes to simulate real-world scenarios. Specifically, we utilize two widely used mobile-captured datasets: On-the-go Ren2024NeRF and RobustNeRF sabour2023robustnerf , containing multiple distractor-scenes in outdoor and indoor environments, respectively. For fair comparison in real data, we train all models on On-the-go outdoor-scenes except Arcdetriomphe and Mountain, which, along with the RobustNeRF indoor-scenes, serve as test scenes. In addition to real data, we also construct synthetic datasets for additional verification. Specifically, we randomly insert distractor from COCO lin2014microsoft into Re10K zhou2018stereo and ACID liu2021infinite data, with details described in Appendix.
@@ 5.1.1 Real and Synthetic Distractor Datasets
5.1.2 Training and Evaluation Setting
In all experiments, we set the number of references $N$=4 and the size of scene-images pool $K$=8, which are discussed in Fig. 10. During all training, query views are randomly selected and reference views are chosen following the nearest-neighbor selection strategy, regardless of ‘clutter’ or ‘extra’. In evaluation phase, we utilize all ‘extra’ images and ‘clear’ images with a stride of eight as query views for On-the-go (Arcdetriomphe and Mountain) and all RobustNeRF scenes. For references in evaluation, we construct the scene-images pool using views closest to the query view, ensuring inclusion of both distractor-containing and few distractor-free data to validate the effectiveness of Reference Scoring mechanism. Note that this setting is only for validation and evaluation purposes. In practical applications, the scene-images pool can be constructed using any adjacent views, independent of the query view and distractor presence. Finally, we compute the average PSNR, SSIM, and LPIPS metrics across all scenes on all query renders. More details are described in Appendix.
@@ 5.2.1 Benchmark
Our Distractor-free Generalizable training and inference paradigms can be seamlessly integrated with existing generalizable 3DGS frameworks. We adopt Mvsplat chen2024mvsplat as our baseline model. Extensive comparisons are conducted against existing works trained under same settings and distractor datasets, including: (1) retraining existing generalization methods chen2024mvsplat ; charatan2024pixelsplat , and (2) retraining Mvsplat chen2024mvsplat ‘+’ different mask prediction strategies from distractor-free approaches Ren2024NeRF ; sabour2023robustnerf ; chen2024nerf ; sabour2024spotlesssplats . Although all current distractor-free works are scene-specific, most of them focus on applying distractor masks in loss function (as discussed in Sec. 2.2), so we can directly transfer them to the query rendering loss under our baseline. Specifically, during the training phase, we combine these mask prediction methods from existing works and replace $\mathcal{M}_{Rob}$ in Eq .3. Furthermore, (3) we compare with existing pre-trained generalizable models (on distractor-free data). More details are provided in Appendix.
5.2.2 Quantitative and Qualitative Experiments
@@
Tab. 1, Fig. 5 and Fig. 7 quantitatively and qualitatively compares DGGS-TR (only TRaining in Sec. 4.1 without two stages inference in Sec. 4.2) and DGGS with existing methods. The results are analyzed from three aspects: re-training models, pre-training models and models after fine-tuning.
For Re-train Model:
@@
Tab. 1 and Fig. 5 demonstrate that our training paradigm significantly enhances training robustness, outperforming existing generalizable 3DGS algorithms, PSNR: 21.02 vs. 15.45, under same training conditions. Compared to introducing existing scene-specific distractor-free approaches, where overaggressive distractor prediction degrades reconstruction fidelity, DGGS-TR exhibits enhanced reconstruction detail and 3D consistency, PSNR: 21.02 vs. 19.29, through precise distractor prediction, which is further validated in Fig. 6. Similarly, the qualitative results in Fig. 5 show training instability and significant performance drops, even without distractors in some test references. It confirms that distractor presence severely impairs model learning of geometric 3D consistency, aligning with our motivation. Additionally, experiments also show that DGGS possesses cross-scene feedforward inference capabilities that existing distractor-free methods lack.
For Pre-train Model:
@@
Tab. 1 reports comparisons between DGGS-TR, DGGS and existing pre-train generalizable models. Compared with pre-train models on distractor-free data, DGGS-TR exhibits superior performance even with training on limited distractor scenes, mainly due to mitigating the inference scene distractor impacts and domain gap. Fig. 7 also illustrates similar findings: DGGS-TR effectively attenuates partial distractor effects in the 3D inconsistency regions. Furthermore, after introducing our inference paradigm, DGGS demonstrates ‘pseudo-completion’ and artifact removal capabilities, which benefit from the Reference Selection mechanism that can choose references with less distractor as well as disparity and the Distractor Pruning during inference process as in Fig. 10. To further validate effectiveness, we compare DGGS’s results on synthetic data, which maintains the same data domain as the pre-trained MVSplat* except for inserted distractor. In Fig. 10, DGGS shows better resistance to distractor, effectively mitigating artifact. More results are shown in Appendix.
After Fine-tuning:
@@
To further demonstrate DGGS’s capability, we conduct fine-tuning experiments on the different pre-trained models (including Mvsplat*, (Mvsplat+SLS)* and DGGS-TR* corresponding Line 2, 8, and 9 in Tab. 1) and training strategies (including FT, SLS-FT and DGGS-FT corresponding Mvsplat chen2024mvsplat , SLS sabour2024spotlesssplats and Ours). We fine-tune using the ‘clutter’ data and evaluate on the ‘extra’ data in test scenes. For fairness in comparison, our inference framework is not employed. In Tab. 5, DGGS-TR demonstrates significant improvements over existing models and strategies, attributed to our reliable distractor prediction. More qualitative results are shown in Appendix.
Efficiency:
@@
Tab. 5 shows the efficiency comparison. Due to the segmentation model and two-stage inference, DGGS has a slight decrease in efficiency compared to chen2024mvsplat , but mitigates distractor and obtains masks. It can be boosted by reducing segmentation resolution or lighter segmentation models.
@@ 5.3.1 Ablation on Training Framework
The upper section of Tab. 5 and Fig. 3 show the effectiveness of each component in DGGS training paradigm. The Reference-based Mask Prediction combined with Mask Refinement successfully mitigates the tendency of over-predicting targets as distractor that occurred in the original $\mathcal{M}_{Rob}$. Within the Mask Refinement module, Auxiliary Loss demonstrates remarkable performance, while Mask Decoupling and Entity Segmentation contribute substantial improvements to the overall framework. Furthermore, our analysis in Fig. 6 reveals that DGGS-TR achieves scene-agnostic mask prediction capabilities, with feedforward inference results superior to single-scene trained models chen2024nerf . Note that, for fair evaluation, all masks are predicted under DGGS-TR based same references, without the involvement of our inference framework.

### v3

@@ 4.1.2 Mask Refinement
Finally, we decouple $\mathcal{M}_{D}$ from $\mathcal{M}_{Q}$ and recombine them after introducing the segmentation model (Qi et al., 2022) and refining the distractor mask. The final refined mask, termed $\mathcal{M}$ in Fig. 3, substitutes $\mathcal{M}_{Rob}$ in Eq. 3 to mitigate distractor effects during training. Furthermore, we observe that if only Mask Refinement is used, without leveraging reference and our observation, the misclassification remains severe in Fig. 3, which verifies the importance of reference filter.
@@
$\hskip-19.91684pt\mathcal{L}_{A}=\sum_{N}^{i=1}\mathcal{W}_{T\rightarrow i}(1-\mathcal{{M}})\odot\mathcal{M}^{En}_{Ref_{i}}\odot\left\|\mathbf{I}_{i}-\mathcal{G}\left(\mathbf{P}_{i}\right)\right\|_{2}^{2}.$ (7)
@@ 4.2 Distractor-free Generalizable Inference
Despite improvements in training robustness and mask prediction, DGGS’s inference faces two key limitations: (1) insufficient references compromise reliable reconstruction of occluded and unseen regions; (2) persistent distractor in references inevitably appear as artifacts in synthesized novel views for feed-forward inference paradigm (encoder-decoder) in Eq. 2. To address these challenges, we propose a two-stage Distractor-free Generalizable Inference framework in Fig. 4. The first stage employs a Reference Scoring mechanism to score candidate references from the images pool, facilitating the selected references with minimal distractor and disparity. The second stage uses a Distractor Pruning module to suppress the remaining artifacts.
Reference Scoring Mechanism.
@@
The objective of the first inference stage is to select references with minimal distractor and disparity among the predefined scene-images pool, adjacent $K$ views ($K\geq N$) sampled in the test scenes. Therefore, we propose a Reference Scoring mechanism based on the pre-trained DGGS in Sec. 4.1. Specifically, it first samples $N$ adjacent references from the scene-images pool $\{\mathbf{I}\}^{K}_{P}$ for coarse 3DGS prediction. We then designate unselected views from $\{\mathbf{I}\}^{K}_{P}$ as query for mask prediction $\{\mathcal{M}\}^{K-N}$, while the distractor masks of the chosen reference views are $\{\mathcal{M}^{En}_{Ref}\}^{N}$. All image masks in pool are scored by Eq. 9, in which top $N$ images are selected in next stage,
$\hskip-11.38092pt\{\mathbf{I}_{i}\}^{N}=\{\mathbf{I}_{i}\}^{K}_{P}\mid i\in\arg\max_{N}\left\{\mathcal{S}\left(\{\mathcal{M}\}^{K-N};\{\mathcal{M}^{En}_{Ref}\}^{N}\right)\right\}.$ (9)
@@
where $\mathcal{S}$ is the pixel-wise summation for each mask. In practice, besides distractor size, the extrinsics of candidate images are also crucial reference scoring factors due to disparity. Thanks to the discussion of disparity-induced error masks in Sec. 4.1.2, we can directly utilize the count of positive pixels in $\mathcal{M}$ as the single criterion, which selects references that provide better coverage of query view, as shown in Fig. 7 9. In the second stage, we employ top-ranked images as references for fine reconstruction, effectively reselecting the candidate images in the pool without increasing GPU memory. Although this approach successfully handles distractor-heavy references, it comes at the cost of decreased rendering efficiency. Optionally, we mitigate this by halving the image resolution in the first stage.
Distractor Pruning.
@@
Although cleaner references are selected, obtaining $N$ distractor-free images in the wild is virtually impossible. These residual distractor propagate via the gaussian encoding-decoding process in Eq. 2, manifesting as phantom splats in rendered query view, as shown in Fig. 9. Therefore, we propose a Distractor Pruning protocol in the second inference stage, which is readily implementable given the reference distractor masks. Instead of direct masking on the references, which affects one-to-one mapping between pixels and gaussian primitives in Eq. 2, we selectively prune gaussian primitives within the 3D space by directly removing decoded attributes in distractor regions while preserving the remaining components. However, when references exhibit a large amount of commonly occluded regions, the pruning strategy induces white speckle artifacts. Consequently, based on projected masks, DGGS implements the pruning strategy only in scenarios where it is not considered a common occluded region for all references, which is discussed in Sec. 6.
@@ 5 Experiments
Pixelsplat* (2024 CVPR) 18.65 0.673 0.254 17.98 0.557 0.364 20.10 0.704 0.279 Pre-Train
Mvsplat* (2024 ECCV) 18.88 0.670 0.225 18.24 0.586 0.301 20.03 0.722 0.255 on Re10K
Pixelsplat (2024 CVPR) 15.49 0.378 0.531 16.34 0.331 0.492 16.02 0.422 0.511
Mvsplat (2024 ECCV) 15.05 0.412 0.391 16.17 0.509 0.381 15.45 0.515 0.426
+RobustNeRF (2023 CVPR) 16.17 0.463 0.382 16.46 0.470 0.411 17.11 0.534 0.400 Re-Train on
+On-the-go (2024 CVPR) 14.73 0.366 0.522 15.05 0.440 0.472 15.44 0.476 0.526 Distractor-
+NeRF-HuGS (2024 CVPR) 18.21 0.694 0.266 18.33 0.640 0.299 19.18 0.700 0.283 Datasets
+HybridGS (CVPR 2025) 17.16 0.540 0.369 16.37 0.517 0.375 17.82 0.556 0.388
+SLS (Arxiv 2024) 18.11 0.695 0.270 18.84 0.662 0.282 19.29 0.709 0.286
@@
+++Mask Refinement (DGGS-TR) 21.02 0.738 0.242
@@
5.1 Synthetic Dataset
To train and evaluate DGGS, beyond the real-world datasets On-the-go (Ren et al., 2024) and RobustNeRF Dataset (Sabour et al., 2023), we construct a synthetic dataset to augment the number of distractor scenes based on Re-10K and ACID, with details provided in AppendixA.
@@ 5.2 Comparative Experiments
Benchmark: Our Distractor-free Generalizable training and inference paradigms can be seamlessly integrated with existing generalizable 3DGS frameworks. We adopt Mvsplat (Chen et al., 2024b) as our baseline model. Extensive comparisons are conducted against existing works trained under same settings and distractor datasets, including: (1) retraining existing generalization methods (Chen et al., 2024b; Charatan et al., 2024), and (2) retraining Mvsplat (Chen et al., 2024b) ‘+’ different mask prediction strategies from distractor-free approaches (Ren et al., 2024; Sabour et al., 2023; Chen et al., 2024a; Sabour et al., 2024). Although all current distractor-free works are scene-specific, most of them focus on applying distractor masks in loss function (as discussed in Sec. 2.2), so we can directly transfer them to the query rendering loss under our baseline. Specifically, during the training phase, we combine these mask prediction methods from existing works and replace $\mathcal{M}_{Rob}$ in Eq .3. Furthermore, (3) we compare with existing pre-trained generalizable models (distractor-free).
@@
Quantitative and Qualitative Experiments: Tab. 1, Fig. 5 and Fig. 7 quantitatively and qualitatively compares DGGS-TR (only TRaining in Sec. 4.1 without two stages inference in Sec. 4.2) and DGGS with existing methods. The results are analyzed from three aspects: re-training models, pre-training models and models after fine-tuning.
@@
For Re-train Model: Tab. 1 and Fig. 5 demonstrate that our training paradigm significantly enhances training robustness, outperforming existing generalizable 3DGS algorithms, PSNR: 21.02 vs. 15.45, under same training conditions. Compared to introducing existing scene-specific distractor-free approaches, where overaggressive distractor prediction degrades reconstruction fidelity, DGGS-TR exhibits enhanced reconstruction detail and 3D consistency, PSNR: 21.02 vs. 19.29, through precise distractor prediction, which is further validated in Fig. 6. Similarly, the qualitative results in Fig. 5 show training instability and significant performance drops, even without distractors in some test references. It confirms that distractor presence severely impairs model learning of geometric 3D consistency, aligning with our motivation. Additionally, experiments also show that DGGS possesses cross-scene feed-forward inference capabilities that existing distractor-free methods lack.
@@
For Pre-train Model: Tab. 1 reports comparisons between DGGS-TR, DGGS and existing pre-train generalizable models. Compared with pre-train models on distractor-free data, DGGS-TR exhibits superior performance even with training on limited distractor scenes, mainly due to mitigating the inference scene distractor impacts and domain gap. Fig. 7 also illustrates similar findings: DGGS-TR effectively attenuates partial distractor effects in the 3D inconsistency regions. Furthermore, after introducing our inference paradigm, DGGS demonstrates ‘pseudo-completion’ and artifact removal capabilities, which benefit from the Reference Selection mechanism that can choose references with less distractor as well as disparity and the Distractor Pruning during inference process as in Fig. 9. To further validate effectiveness, we compare DGGS’s results on synthetic data, which maintains the same data domain as the pre-trained MVSplat* except for inserted distractor. In Fig. 9, DGGS shows better resistance to distractor, effectively mitigating artifact. More results are shown in Appendix.
@@
After Fine-tuning: To further demonstrate DGGS’s capability, we conduct fine-tuning experiments on the different pre-trained models (including Mvsplat*, (Mvsplat+SLS)* and DGGS-TR* corresponding Line 2, 8, and 9 in Tab. 1) and training strategies (including FT, SLS-FT and DGGS-FT corresponding Mvsplat (Chen et al., 2024b), SLS (Sabour et al., 2024) and Ours). We fine-tune using the ‘clutter’ data and evaluate on the ‘extra’ data in test scenes. For fairness in comparison, our inference framework is not employed. In Tab. 4, DGGS-TR demonstrates significant improvements over existing models and strategies, attributed to our reliable distractor prediction. Furthermore, we find that DGGS’s single-scene fine-tuning performance is even better than SLS’s single-scene training (metrics reported by SLS (Sabour et al., 2024)). This benefits from DGGS’s pretraining on large-scale datasets and its reference-based strategy for distractor pridiction.
@@
Efficiency: Tab. 4 shows the efficiency comparison. Due to the segmentation model and two-stage inference, DGGS has a slight decrease in efficiency compared to (Chen et al., 2024b), but mitigates distractor and obtains masks. It can be boosted by reducing segmentation resolution.
Re-rendered Reference Accuracy: One of the core insights of DGGS is to leverage the more stable reference re-render as an additional filter to assist query distractor mask prediction. Most existing works exhibit this trend, such as Mvsplat* (Chen et al., 2024b). To verify the stability of re-rendering during the generalizable 3DGS entire training process, we compare the performance of query and re-rendering reference views in Tab. 6. From the experimental results on DGGS, it can be easily observed that from the beginning of DGGS training, reference re-rendering exhibits more stable performance relative to the query view. This characteristic allows reference re-rendering to effectively guide distractor mask prediction for the query view. Furthermore, we simultaneously observe that as the training process progresses, the performance gap between the two gradually narrows.
@@ 5.3.1 Ablation on Training Framework
The upper section of Tab. 4 and Fig. 3 show the effectiveness of each component in DGGS training paradigm. The Reference-based Mask Prediction combined with Mask Refinement successfully mitigates the tendency of over-predicting targets as distractor that occurred in the original $\mathcal{M}_{Rob}$. Within the Mask Refinement module, Auxiliary Loss demonstrates remarkable performance, while Mask Decoupling and Entity Segmentation contribute substantial improvements to the overall framework. Furthermore, our analysis in Fig. 6 reveals that DGGS-TR achieves scene-agnostic mask prediction capabilities, with feed-forward inference results superior to single-scene trained models (Chen et al., 2024a). Note that, for fair evaluation, all masks are predicted under DGGS-TR based same references, without the involvement of our inference framework.
Table 5: Ablation Study on Segmentation Module under RobustNeRF dataset.
[…]
DGGS-TR (w/o Segmentation) 20.67 0.730 0.251
DGGS-TR (SAM2) 21.07 0.740 0.240
DGGS-TR (w/ Segmentation) 21.02 0.738 0.242
DGGS (w/o Segmentation) 21.41 0.749 0.240
DGGS (SAM2) 21.77 0.758 0.236
DGGS (w/ Segmentation) 21.74 0.758 0.237
Table 6: Query v.s. Re-rendering Reference Analysis comparing PSNR at different training iterations.
Method Train Stage Query Ref Re-render
Trained Mvsplat* Done 26.02 29.71
DGGS-TR 200 12.13 17.82
DGGS-TR 1000 16.36 20.38
DGGS-TR 10000 22.11 25.65
Trained DGGS-TR Done 26.51 28.02